OpenStudy (anonymous):

Sean bought a bag of mixed candies. In the bag there were 25 Smarties and Reese's Peanut Butter Cups. Smarties cost $0.15 each and Reese's Cups are $0.75 each. If he paid $15.15, how many Smarties and Reese's Cups did he buy? A. Write an expression for the number of Smarties. Write an expression for the number of Reese's Cups. B. Write an algebraic equation using the expressions above. C. Solve your equation. Show all work. D. Tell how many Smarties and Reese's Cups were bought. Can you tell me if my answer is right?

OpenStudy (anonymous):

A. s = the amount of smarties 25 - s = the amount of Reese's B. s X $0.15 + (25 - s) X $0.75 = $15.15 C. s X $0.15 + (25 - s) X $0.75 = $15.15 $0.15s + $0.75(25 - s) = $15.15 $0.15s + $18.75 - $0.75s = $15.15 $18.75 - $0.6s = $15.15 $18.75 - $18.75 - $0.6s = $15.15 - $18.75 -$0.6s = -$3.60 -$0.6s/-$0.6 = -$3.60/-$0.6 s = 6 D. Sean bought six Smarties and 19 Reese's Cups. That's my answer

OpenStudy (dumbcow):

you can always plug it in to check your answer 6+19 = 25, check 6(.15) + 19(.75) = 15.15 , check